OpenStudy (anonymous):

Solve the following system of equations using the elimination method. If the system is inconsistent, type INCONSISTENT in both boxes. If the system is dependent, type DEPENDENT in both boxes. 4x+7y=-1 3x+3y=-3

OpenStudy (whpalmer4):

to do elimination, you want to multiply one or both equations by numbers such that you'll end up with either 2 coefficients with an equal value, or 2 coefficients with equal values but opposite signs. At that point, you respectively subtract or add the two equations, the two matching terms vanish, and you are left with an equation in 1 fewer variable. Solve that, then use the value found there in any of the original equations to find the other variable's value. If you end up with a situation like 0=0, that means you have a dependent system. If you end up with a situation like 0=1, that means you have an inconsistent system.

OpenStudy (anonymous):

Let me get you started... 4x + 7y = - 1 -->(3)4x + 7y = - 1 3x + 3y = - 3 -->(-4)3x + 3y = - 3 ---------------- 12x + 21y = - 3 (result of multiplying by 3) -12x - 12y = 12 (result of multiplying by -4) ----------------add 0 + 9y = 9 9y = 9 y = 1 now just sub 1 in for y in either of the original equations to find x. As you can see, I multiplied the equations to eliminate the x's , thus solving for y.